OpenStudy (anonymous):

The population of a city has been continuously declining over the past 15 years at a rate of 8%. Use the formula A = Pe^rt, where P is the initial population, t is the time in years, to find the population 15 years ago if the population is currently 37,469 people. Select one: a. 12,401 b. 102,401 c. 124,401 d. 186,401

OpenStudy (anonymous):

i think C.

OpenStudy (anonymous):

i was thinking B

OpenStudy (anonymous):

why do you think B ?

OpenStudy (yanasidlinskiy):

None of the answers are close. You may just want to take a guess.

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

thats what i was thinking . theyre not close but i would guess C. Do what you want to do .

OpenStudy (acxbox22):

these answer choices are not correct

OpenStudy (yanasidlinskiy):

Look: P = A / (0.92 ^ 15) = 37469 / 0.2863 = 130877

OpenStudy (anonymous):

thats what i got... thats why i would just go with the highest number.

OpenStudy (yanasidlinskiy):

Just take a guess. And yea. Just go with the highest I guess:)

OpenStudy (anonymous):

well the closest would be C

OpenStudy (yanasidlinskiy):

Yep!

OpenStudy (acxbox22):

\[x(0.92)^{15}=37469\] \[x(0.28629740407)=37469\] \[x \approx 130874.39844 \]

OpenStudy (acxbox22):

according to my answer C is the most reasonable choice